NH4NO3(s) → N2O(g) + 2H2O(l)


In a certain experiment, 156 g of ammonium nitrate is decomposed. Find the mass of each of the products formed.

1
Expert's answer
2022-05-20T04:27:03-0400

"M(NH_4NO_3)=14+4x1+14+16x3=80 g\/mole"

"M(N_2O)=14x2+16=44 g\/mole"

"M(H_2O)=2x1+16=18 g\/mole"


"n(NH_4NO_3)=156\/80=1.95 mole"

"n(N_2O)=n(NH_4NO_3)=1.95 mole"

"n(H_2O)=2n(NH_4NO_3)=1.95x2=3.9 mole"


"m(N_2O)=1.95x44=85.8 g"

"m(H_2O)=3.9x18=70.2 g"